# A self-supervised framework for learning whole slide representations

### v1

@@ C.3 Extended ablation results
We present additional ablation experiments varying cropping size range in Figure 9. Cropping size is a hyperparameter that allows us to adjust the amount of mutual information between views, i.e. larger crop size $\rightarrow$ higher probability and area of overlapping regions between views. Our ablation study shows that for SimCLR training on our H&E benchmark, $[100,400]$ is the optimal cropping size range, achieving a MCA of 76.5. These results correspond to the hypothesis in (Tian et al., 2020), where an optimal amount of mutual information between views achieves the best model performance.
@@
Figure 9: Cropping size range ablation for H&E. Cropping size range is on the x-axis, where numbers in the bracket correspond to the min and max cropping size. None, no cropping applied, Shaded area, standard deviation across three different random seeds.
C.4 Extended attention heatmaps
@@
Figure 10: Extended attention heatmap for H&E whole slides. The attention map shows the S3L-trained whole slide transformer can differentiate between different morphologies like dense tumors, blood, and different artifacts on the gigapixel WSIs. Visualization shows high attention to varying degrees of tumor infiltration, and low attention to low cellularity, blood, necrotic regions, and tissue processing artifacts. Oligo, oligodendroglioma, Astro, astrocytoma, GBM, glioblastoma.

### v2

@@ C.5.3 SPT transformation parameters
We varied cropping size range in Figure 9b. Cropping also allows us to adjust the amount of MI between views, where a larger crop size leads to a higher probability of overlapping regions between views. Our ablation study shows that for SimCLR training on our H&E glioma benchmark, $[100,400]$ is the optimal cropping size range, achieving a MCA of 76.5. These results correspond to the hypothesis in [44], where an optimal amount of MI between views achieves the best performance.
@@
Figure 9: SPT transformation parameter ablation for H&E glioma molecular classification. a) Max token limit ablation. SCM, splitting-cropping-masking transformation. b) Cropping size range ablation. Cropping size range [min, max cropping size] is on the x-axis. None represent no cropping applied. The shaded area represents the standard deviation across three different random seeds.
C.5.4 SPT model size
We performed additional ablation studies with different SPT model sizes, and the results are reported in Table 13. The six- and two-layer transformers have the best performance for H&E glioma and SRH CNS, respectively. Smaller models performed better for SRH CNS, likely because of a smaller slide size and a relatively more uniform image. All models ranging from two to eight-layer transformers outperformed previous best self-supervised models.
SRH H&E Glioma
# Layers # Parameters MCA F1 AUC MCA F1 AUC
2 6.7 M 83.5 (0.1) 83.8 (0.3) 95.4 (0.4) 74.9 (0.9) 74.8 (1.0) 91.6 (0.5)
4 13 M 82.2 (0.9) 83.2 (1.6) 95.4 (0.2) 75.5 (0.8) 75.3 (0.5) 91.1 (0.4)
6 19.3 M 82.3 (0.5) 82.3 (0.3) 94.7 (0.1) 76.5 (0.5) 76.1 (0.4) 90.9 (0.4)
8 25.6 M 81.4 (0.2) 82.5 (0.3) 95.2 (0.2) 75.2 (1.2) 75.2 (1.1) 91.7 (0.9)
Table 13: SPT ablation studies on model size. Standard deviations are in (parentheses).
C.6 Extended attention heatmaps
@@
Figure 10: Extended attention heatmap for H&E whole slides. The attention map shows the ssSPT-trained whole slide transformer can differentiate between different morphologies like dense tumors, blood, and different artifacts on the gigapixel WSIs. Visualization shows high attention to varying degrees of tumor infiltration, and low attention to low cellularity, blood, necrotic regions, and tissue processing artifacts. Oligo, oligodendroglioma, Astro, astrocytoma, GBM, glioblastoma.